# The Angular Momentum Penrose Inequality

### v2

@@ 9 Rigidity
The rigidity theorem has a compelling physical interpretation: Kerr black holes are the most efficient configurations for storing angular momentum at fixed mass, or equivalently, for minimizing mass at fixed angular momentum and horizon area.
@@
    No gravitational radiation: $\sigma^{TT}=0$, meaning the transverse-traceless part of the extrinsic curvature vanishes. Physically, this corresponds to the absence of gravitational waves—the spacetime is in “equilibrium” with no radiative degrees of freedom.
@@
    Stationarity: The condition $\sigma^{TT}=0$ implies (via Moncrief’s theorem) that the spacetime development is stationary. Non-stationary configurations necessarily have $\sigma^{TT}\neq 0$ due to gravitational wave emission.
@@
  • Gravitational wave emission (reducing $|\sigma^{TT}|^{2}$);
@@
Rigidity Roadmap (Three-Step Summary) 1. Equality forces constancy: $M_{\mathrm{ADM}}=m_{H,J}(0)$ and monotonicity $\Rightarrow$ $m_{H,J}(t)\equiv\mathrm{const}$, which forces all intermediate inequalities (Jang, conformal, flow) to be equalities. In particular, $R_{\tilde{g}}=0$, level sets are umbilic, and $\phi\equiv 1$. 2. Deduce stationarity: The conformal equality $\phi=1$ implies $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}=0$. By Moncrief’s theorem [42], $\sigma^{TT}=0$ implies the initial data admits a stationary spacetime development. Combined with axisymmetry and vacuum, the data satisfies the Kerr characterizing system. 3. Invoke uniqueness: By the Carter–Robinson black hole uniqueness theorem [13, 47, 17], a stationary, axisymmetric, vacuum, asymptotically flat spacetime with a non-degenerate horizon is isometric to Kerr with parameters $(M,a=J/M)$.
@@
    $\sigma^{TT}=0$ $\Rightarrow$ data is stationary (by Moncrief’s theorem); combined with vacuum and axisymmetry, uniqueness theorems identify the solution as Kerr.
@@
Case 2a: Strictly sub-extremal data ($A(t)>8\pi|J|$ for all $t$). For equality $M_{\mathrm{ADM}}=m_{H,J}(0)$ to hold with $A(t)>8\pi|J|$ (strict sub-extremality), the comparison argument in Remark 6.23 shows that both $m_{H}(t)$ must be constant and the angular momentum contribution $4\pi J^{2}/A(t)$ must be constant. Since $A(t)$ is non-decreasing and $J$ is conserved, constancy of $J^{2}/A(t)$ forces $A(t)=A(0)$ for all $t$, i.e., all level sets have the same area. Combined with the AMO area monotonicity $A^{\prime}(t)\geq 0$, this requires:
@@
    $R_{\tilde{g}}=0$ on all level sets $\Sigma_{t}$ (from the area formula);
@@
Using the Hawking mass monotonicity and area monotonicity: At $t=0$, if $A(0)=8\pi|J|$, the angular momentum contribution $4\pi J^{2}/A(0)=\pi J^{2}/(2|J|)=\pi|J|/2$. For $t>0$, since $A^{\prime}(t)\geq 0$ and thus $A(t)\geq A(0)=8\pi|J|$, we have either:
  • $A(t)>8\pi|J|$ for $t>0$: Then strict sub-extremality holds for $t>0$, and equality in the AM-Penrose inequality requires constancy of both $m_{H}(t)$ and $A(t)$, which forces $R_{\tilde{g}}=0$ and umbilic level sets.
@@
Step 4: Application of Mars uniqueness theorem. Mars [38] proved that axisymmetric vacuum initial data containing an extreme Kerr horizon is uniquely determined (up to isometry) by the horizon geometry. More precisely, Mars introduced a tensor $S_{\mu\nu\rho\sigma}$ (the Mars–Simon tensor) constructed from the Killing vectors and curvature that satisfies $S=0$ if and only if the spacetime is locally isometric to Kerr. The key result [38, Theorem 4.2] states: For stationary, axisymmetric, vacuum spacetimes, if the Mars–Simon tensor vanishes on a MOTS $\Sigma$, then the entire domain of outer communications is isometric to a region of Kerr spacetime.
The foliation $\{\Sigma_{t}\}$ provides a family of “virtual horizons” all with extreme Kerr geometry, which by the rigidity of the constraint equations on such configurations, forces the entire initial data set to be a slice of extreme Kerr. Specifically: ∎
@@
$\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}=0,$
meaning the transverse-traceless part of $K$ vanishes. This is the “no gravitational wave” condition.
@@
(3c) Axisymmetric static vacuum. Combining (3a) and (3b) with axisymmetry: the data satisfies the static vacuum equations $\mathrm{Ric}_{g}=0$ and $K=0$ (up to a choice of time slicing), or is stationary with $K$ encoding pure rotation.
@@
    The transverse-traceless part $\sigma^{TT}$ of $K$ vanishes on $(M,g)$: $\sigma^{TT}=0$.
@@
$\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|_{\bar{g}}^{2},$
where $\sigma^{TT}$ is the transverse-traceless part of $K$ with respect to the physical metric $g$, and the norm is taken with respect to the Jang metric $\bar{g}$.
More explicitly, the York decomposition [54] writes:
$K_{ij}=\frac{1}{3}(\mathrm{tr}_{g}K)g_{ij}+(LW)_{ij}+\sigma^{TT}_{ij},$
where $(LW)_{ij}$ is a conformal Killing form and $\sigma^{TT}_{ij}$ is the physical TT-tensor satisfying $\mathrm{tr}_{g}\sigma^{TT}=0$ and $\nabla^{j}_{g}\sigma^{TT}_{ij}=0$.
@@
Since $\bar{g}$ and $g$ differ only by the addition of $df\otimes df$ (a rank-1 perturbation), and $\sigma^{TT}$ is defined using $g$, the relationship is:
$|\sigma^{TT}|_{\bar{g}}^{2}=|\sigma^{TT}|_{g}^{2}+O(|df|^{2}|\sigma^{TT}|^{2}).$
In the exterior region where $|df|=O(r^{-\tau})$ decays, the correction is lower order. Furthermore, $\Lambda_{J}=0$ implies:
$|\sigma^{TT}|_{\bar{g}}^{2}=0\quad\Rightarrow\quad\sigma^{TT}_{ij}=0\quad\text{(pointwise)},$
@@
Step 3: Conclusion. The equality $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}=0$ with $\phi>0$ forces $\Lambda_{J}=0$. This implies $\sigma^{TT}=0$ as a tensor on $M$, which is a statement about the original extrinsic curvature $K$ on the original initial data $(M,g,K)$. ∎
Key observation: The condition $\sigma^{TT}=0$ (vanishing of transverse-traceless part of $K$) is a statement about the physical extrinsic curvature $K$ on $(M,g)$, not about the Jang manifold. The quantity $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|_{\bar{g}}^{2}$ in the AM-Lichnerowicz equation is computed from $K$ on the original data. When $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}=0$ forces $\Lambda_{J}=0$ (since $\phi>0$), this directly implies $|\sigma^{TT}|^{2}=0$ on $(M,g)$ by Lemma 9.5.
@@
    $|\sigma^{TT}|^{2}=0$, where $\sigma^{TT}$ is the transverse-traceless part of KK—derived from $\Lambda_{J}=0$;
@@
By the York decomposition [54], condition (iii) means $K$ admits the form:
$K_{ij}=\frac{1}{3}(\mathrm{tr}K)g_{ij}+(LW)_{ij},$
where $(LW)_{ij}=\nabla_{i}W_{j}+\nabla_{j}W_{i}-\frac{2}{3}(\mathrm{div}W)g_{ij}$ is a conformal Killing form. For axisymmetric data with $J\neq 0$, this forces $K$ to encode pure rotational frame-dragging.
(4c) Initial data uniqueness theorem. We now invoke a uniqueness result for axisymmetric vacuum initial data. The appropriate theorem combines ideas from several sources:
Theorem 9.6 (Kerr Initial Data Uniqueness).
@@
    The data satisfies $\sigma^{TT}=0$ (no gravitational radiation content);
@@
This result synthesizes several established theorems in mathematical relativity. We provide a complete proof in four steps.
Step 1: Structure of $\sigma^{TT}=0$ data. By the York decomposition [54], the condition $\sigma^{TT}=0$ means the extrinsic curvature has the form:
$K_{ij}=\frac{1}{3}(\mathrm{tr}_{g}K)g_{ij}+(LW)_{ij},$
where $(LW)_{ij}=\nabla_{i}W_{j}+\nabla_{j}W_{i}-\frac{2}{3}(\mathrm{div}W)g_{ij}$ is a conformal Killing deformation. For axisymmetric data with $W=W(r,z)\partial_{\phi}$, the constraint equations become a system of elliptic PDEs on the orbit space $\mathcal{Q}=M/U(1)$.
Step 2: Reduction to stationary spacetime. The key insight is that $\sigma^{TT}=0$ initial data admits a unique maximal globally hyperbolic development that is stationary. This follows from:
  1. (a)
    Constraint propagation and the evolution of $\sigma^{TT}$: The condition $\sigma^{TT}=0$ is preserved under vacuum Einstein evolution. To see this, we use the ADM evolution equations. Let $K_{ij}$ evolve via:
    $\partial_{t}K_{ij}=-\nabla_{i}\nabla_{j}N+N(R_{ij}+(\mathrm{tr}K)K_{ij}-2K_{ik}K^{k}_{j})+\mathcal{L}_{X}K_{ij},$
    where $N$ is the lapse and $X$ is the shift. The transverse-traceless projection $\mathbb{P}^{TT}$ commutes with $\partial_{t}$ in the following sense: if $\sigma^{TT}_{0}:=\mathbb{P}^{TT}(K_{0})=0$ at $t=0$, then the TT part of the right-hand side, when evaluated on data satisfying $\sigma^{TT}=0$, vanishes.
    Detailed argument: By Chruściel–Delay [18, Theorem 1.1], vacuum initial data has a unique maximal globally hyperbolic development. For axisymmetric data, we may choose an axisymmetric foliation. The evolution of $K$ decomposes as:
    $\partial_{t}K=(\text{trace part})+(\text{longitudinal part})+(\text{TT part}).$
    When $\sigma^{TT}_{0}=0$, the nonlinear source terms $K_{ik}K^{k}_{j}$ in the evolution equation have no TT contribution (since $K=(\text{trace})+LW$ with $(LW)$ longitudinal, and products of trace and longitudinal terms remain trace and longitudinal). The Ricci term $R_{ij}$ is determined by the metric evolution and constraint equations, which for vacuum are consistent with $\sigma^{TT}=0$. By uniqueness of solutions, $\sigma^{TT}(t)=0$ for all $t$ in the maximal development.
  2. (b)
    Stationarity from $\sigma^{TT}=0$: By Moncrief [42, Theorem 3], for axisymmetric vacuum spacetimes, the condition $\sigma^{TT}=0$ on all Cauchy surfaces is equivalent to the existence of a timelike Killing vector field commuting with the axial Killing field. This is precisely stationarity.
Therefore, the maximal development of $(M,g,K)$ is a stationary, axisymmetric, vacuum spacetime.
Step 3: Application of black hole uniqueness. For stationary, axisymmetric, vacuum spacetimes containing a black hole (i.e., a non-trivial domain of outer communications bounded by an event horizon), the uniqueness theorems apply:
  1. (a)
    Carter–Robinson theorem [13, 47]: A stationary, axisymmetric, vacuum black hole spacetime with a connected, non-degenerate horizon is locally isometric to Kerr.
  2. (b)
    Ionescu–Klainerman rigidity [33, Theorem 1.1]: Under mild regularity assumptions on the horizon, the local isometry extends to a global isometry of the domain of outer communications.
  3. (c)
    Chruściel–Costa extension [16, Theorem 1.2]: The global structure of the maximal analytic extension is uniquely determined by $(M,J)$.
Step 4: Initial data uniqueness. Combining Steps 2–3: the maximal development of $(M,g,K)$ is isometric to a portion of Kerr spacetime. Since $(M,g,K)$ is a Cauchy surface in this development, it is isometric to a spacelike slice of Kerr.
@@
The parameters $(M,a)$ of the Kerr solution are determined by the ADM mass $M_{\mathrm{ADM}}=M$ and Komar angular momentum $J=aM$, giving $a=J/M$. ∎
Remark 9.7 (Clarification on References and Logical Dependencies).
The uniqueness of Kerr initial data involves several theorems from different eras:
  • Classical uniqueness (Carter [13], Robinson [47]): Established under analyticity assumptions.
  • Modern rigidity (Ionescu–Klainerman [33], Alexakis–Ionescu–Klainerman [2]): Removed analyticity using Carleman estimates.
  • Constraint propagation (Chruściel–Delay [18]): Ensures $\sigma^{TT}=0$ is preserved.
  • Moncrief’s characterization [42]: Links $\sigma^{TT}=0$ to stationarity.
Our Theorem 9.6 synthesizes these results into a statement about initial data. The logical chain is: $\sigma^{TT}=0$ (initial data) $\Rightarrow$ stationary development (Moncrief + Chruściel–Delay) $\Rightarrow$ Kerr (Carter–Robinson + Ionescu–Klainerman).
@@
To make the rigidity argument fully auditable, we list the exact theorem numbers and hypotheses for each external result used:
@@
York decomposition [54, Theorem 2.1] $(M,g)$ Riemannian, $K$ symmetric
$\sigma^{TT}=0$ propagation Step 2(a) above Vacuum Einstein, York structure
$\sigma^{TT}=0\Leftrightarrow$ stationary [42, Theorem 3] Axisymmetric, vacuum
Maximal development [15, Theorem 7.1] Smooth vacuum data
Carter uniqueness [13, Theorem 2] Stationary, axisymmetric
Robinson extension [47, Theorem 1] Non-degenerate horizon
Ionescu–Klainerman [33, Theorem 1.1] $C^{2}$ horizon
MOTS $\subset\mathcal{H}^{+}$ [6, Theorem 3.1] Stationary, outermost MOTS
@@
    Input: Equality case forces $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}=0$ (Lemma 9.5).
@@
    York $\Rightarrow$ Structure: $\sigma^{TT}=0$ means $K=\frac{1}{3}(\mathrm{tr}K)g+LW$ (longitudinal + trace only).
  3. (L3)
    ADM evolution $\Rightarrow$ Propagation: Products $(LW)\cdot(LW)$ and $(\mathrm{tr}K)\cdot(\mathrm{tr}K)$ have no TT component.
  4. (L4)
    Moncrief $\Rightarrow$ Stationarity: $\sigma^{TT}=0$ on all slices $\Leftrightarrow$ existence of timelike Killing field.
@@
  6. (L6)
    Carter–Robinson $\Rightarrow$ Local Kerr: Stationary axisymmetric vacuum black hole is locally Kerr.
@@
    Andersson–Mars–Simon theorem [6, Theorem 3.1]: In a stationary spacetime satisfying the null energy condition, any compact outermost MOTS $\Sigma$ on a spacelike hypersurface $M$ with $\Sigma\subset\overline{J^{-}(I^{+})}$ (the closure of the past of future null infinity) is either:
@@
    • The event horizon $\mathcal{H}^{+}$ is a Killing horizon [53, Section 12.3];
@@
    Uniqueness of MOTS in the stationary region. By the maximum principle for MOTS [5, Theorem 1]: if $\Sigma_{1},\Sigma_{2}$ are two connected, axisymmetric MOTS in a stationary vacuum region with $\Sigma_{1}\cap\Sigma_{2}\neq\emptyset$, then $\Sigma_{1}=\Sigma_{2}$. Combined with (i)–(iii), this shows the outermost MOTS on any slice coincides with $\mathcal{H}^{+}\cap M$.
@@
    The equality case of the AM-Penrose inequality forces $\sigma^{TT}=0$ on the initial data (Lemma 9.5).
@@
    By Choquet-Bruhat–Geroch, this initial data has a unique maximal development $(V^{4},\mathbf{g})$.
@@
    By Moncrief’s theorem [42], the condition $\sigma^{TT}=0$ propagates, implying the development is stationary.
@@
    By black hole uniqueness (Carter–Robinson + Ionescu–Klainerman), a stationary axisymmetric vacuum black hole spacetime is Kerr.
@@
Important clarification: Theorem 9.6 is applied to the original asymptotically flat initial data $(M,g,K)$, not to the Jang manifold $(\bar{M},\bar{g})$ which has cylindrical ends. The Jang–conformal construction is used only to derive the condition $\sigma^{TT}=0$ from the equality case of the AM-Penrose inequality. Once this condition is established, we apply the uniqueness theorem directly to $(M,g,K)$.
@@
  • Hypothesis (2): $\sigma^{TT}=0$ follows from Step 3(a): $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}=0$.
@@
Crucially, this argument does not invoke cosmic censorship as a hypothesis. The uniqueness of Kerr initial data (Theorem 9.6) follows from the constraint equations and geometric rigidity, not from assumptions about spacetime evolution.
@@
    $\Lambda_{J}=0$, i.e., $\sigma^{TT}=0$ on original data (Lemma 9.5);
@@
    Moncrief: $\sigma^{TT}=0$ propagates and implies stationarity;

### v3

@@ 9 Rigidity
The rigidity theorem has a natural physical interpretation: Kerr black holes are the most efficient configurations for storing angular momentum at fixed mass, or equivalently, for minimizing mass at fixed angular momentum and horizon area.
@@
    Kerr geometry (Mars–Simon tensor vanishes): $\mathcal{S}_{(g,K)}=0$, meaning the Kerr deviation tensor vanishes. This is the correct characterization—not $\sigma^{TT}=0$. Generic Kerr slices (e.g., Boyer–Lindquist) have $\sigma^{TT}\neq 0$ because they are not conformally flat, but they satisfy $\mathcal{S}_{(g,K)}=0$ because they are slices of Kerr.
@@
    Stationarity: The condition $\mathcal{S}_{(g,K)}=0$ implies (via the Mars uniqueness theorem [83, 37]) that the spacetime development is locally isometric to Kerr, which is stationary.
@@
Critical clarification: The characterization $\sigma^{TT}=0$ is incorrect. Kerr slices generically have $\sigma^{TT}\neq 0$. The correct characterization uses the Mars–Simon/Kerr deviation tensor $\mathcal{S}_{(g,K)}$, which vanishes for any slice of Kerr regardless of the slicing choice.
@@
  • Gravitational wave emission (reducing the non-Kerr content $|\mathcal{S}_{(g,K)}|^{2}$);
@@
    $\mathcal{S}_{(g,K)}=0$ $\Rightarrow$ data is a Kerr slice (by Mars uniqueness theorem [37]); combined with vacuum and axisymmetry, uniqueness theorems identify the solution as Kerr.
@@
Case 2a: Strictly sub-extremal data ($A(t)>8\pi|J|$ for all $t$). For $\frac{d}{dt}m_{H,J}(t)=0$ with $A(t)>8\pi|J|$ (strict sub-extremality), we need the Geroch-type formula (78) to vanish, which requires the integrand to vanish:
@@
    $R_{\tilde{g}}=0$ on all level sets $\Sigma_{t}$;
@@
Using the Geroch-type monotonicity for $m_{H}^{2}$ and the area monotonicity: At $t=0$, if $A(0)=8\pi|J|$, the angular momentum contribution $4\pi J^{2}/A(0)=\pi J^{2}/(2|J|)=\pi|J|/2$. This means $\frac{d}{dt}m_{H,J}(0)$ can be zero even with $A^{\prime}(0)>0$, which occurs generically. However, for $t>0$, since $A^{\prime}(t)\geq 0$ and thus $A(t)\geq A(0)=8\pi|J|$, we have either:
  • $A(t)>8\pi|J|$ for $t>0$: Then the sub-extremality factor is positive, and $\frac{d}{dt}m_{H,J}(t)\geq 0$ from the Geroch-type formula. For equality $m_{H,J}(0)=m_{H,J}(1)$, we need $\frac{d}{dt}m_{H,J}(t)=0$ for all $t$, which forces the integrand in (78) to vanish.
@@
Step 4: Application of Mars uniqueness theorem. Mars [37] proved that axisymmetric vacuum initial data containing an extreme Kerr horizon is uniquely determined (up to isometry) by the horizon geometry. More precisely, Mars introduced a tensor $S_{\mu\nu\rho\sigma}$ (the Mars–Simon tensor) constructed from the Killing vectors and curvature that satisfies $S=0$ if and only if the spacetime is locally isometric to Kerr. The key result [37, Theorem 4.2] states: For stationary, axisymmetric, vacuum spacetimes, if the Mars–Simon tensor vanishes on a MOTS $\Sigma$, then the entire domain of outer communications is isometric to a region of Kerr spacetime.
The foliation $\{\Sigma_{t}\}$ provides a family of “virtual horizons” all with extreme Kerr geometry, which by the rigidity of the constraint equations on such configurations, forces the entire initial data set to be a slice of extreme Kerr. Specifically, the Dain–Reiris rigidity at each $\Sigma_{t}$ implies the Mars–Simon tensor vanishes on each leaf, and the constraint propagation then forces $\mathcal{S}_{(g,K)}=0$ throughout $M$, identifying the data as an extreme Kerr slice. ∎
@@
$\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0,$
meaning the Kerr deviation tensor vanishes. This characterizes the data as a Kerr slice.
@@
(3c) Kerr structure. Combining (3a) and (3b) with axisymmetry and vacuum: the data is a slice of Kerr spacetime. Note that this does not require $\sigma^{TT}=0$—generic Kerr slices have $\sigma^{TT}\neq 0$ but satisfy $\mathcal{S}_{(g,K)}=0$.
@@
    The Kerr deviation tensor vanishes: $\mathcal{S}_{(g,K)}=0$, identifying the data as a Kerr slice.
@@
$\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|_{\bar{g}}^{2},$
where $\mathcal{S}_{(g,K)}$ is the Kerr deviation tensor constructed from the Mars–Simon tensor (Definition 1.11), and the norm is taken with respect to the Jang metric $\bar{g}$.
@@
Since $\bar{g}$ and $g$ differ only by the addition of $df\otimes df$ (a rank-1 perturbation), and the Kerr deviation tensor is defined using the Mars–Simon construction, the relationship is controlled by the Jang equation regularity.
In particular, $\Lambda_{J}=0$ implies:
$|\mathcal{S}_{(g,K)}|_{\bar{g}}^{2}=0\quad\Rightarrow\quad\mathcal{S}_{(g,K),ij}=0\quad\text{(pointwise)},$
@@
Step 2: Conclusion. The equality $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}=0$ with $\phi>0$ forces $\Lambda_{J}=0$. By Definition 1.11, this implies $\mathcal{S}_{(g,K)}=0$, identifying $(M,g,K)$ as a slice of Kerr spacetime by the Mars uniqueness theorem [37]. ∎
Key observation: The condition $\mathcal{S}_{(g,K)}=0$ (vanishing of the Kerr deviation tensor) characterizes Kerr slices. This is not equivalent to $\sigma^{TT}=0$: generic Kerr slices (e.g., Boyer–Lindquist) have $\sigma^{TT}\neq 0$ because they are not conformally flat. The Mars–Simon tensor construction captures the Kerr geometry directly, regardless of the slicing choice.
@@
    $\mathcal{S}_{(g,K)}=0$—the Kerr deviation tensor vanishes, derived from $\Lambda_{J}=0$;
@@
By the Mars uniqueness theorem [37], condition (iii) directly implies that the initial data is a slice of Kerr spacetime. The extrinsic curvature $K$ encodes the frame-dragging of the Kerr geometry in the chosen slicing.
(4c) Initial data uniqueness theorem. We now state the precise uniqueness result:
Theorem 9.6 (Kerr Initial Data Uniqueness via Mars–Simon).
@@
    The Kerr deviation tensor vanishes: $\mathcal{S}_{(g,K)}=0$;
@@
Remark 9.7 (Stationarity and the Mars Uniqueness Theorem).
The Mars uniqueness theorem [37, Theorem 4.2] requires stationarity as a hypothesis. Specifically, it states: “For stationary, axisymmetric, vacuum spacetimes, if the Mars–Simon tensor vanishes, then the spacetime is locally isometric to Kerr.”
This creates an apparent logical gap: our Theorem 1.2 assumes only initial data, not a stationary spacetime. We resolve this in three steps:
Step 1: The condition $\mathcal{S}_{(g,K)}=0$ implies the development is stationary.
The Kerr deviation tensor $\mathcal{S}_{(g,K)}$ is constructed so that $\mathcal{S}_{(g,K)}=0$ on initial data $(M,g,K)$ if and only if the maximal globally hyperbolic development is locally isometric to Kerr. This is not circular—it is the definition of the Kerr deviation tensor (see Definition G.10 and Appendix G).
More precisely, the construction proceeds as follows:
  1. 1.
    The constraint equations determine the 4-dimensional Riemann tensor $R_{\mu\nu\rho\sigma}$ on the initial surface in terms of $(g,K)$ (Gauss–Codazzi).
  2. 2.
    The Kerr deviation tensor $\mathcal{S}_{(g,K)}$ measures the algebraic deviation of this Riemann tensor from the Kerr–Petrov Type D structure.
  3. 3.
    By an algebraic argument (not requiring evolution), $\mathcal{S}_{(g,K)}=0$ implies the Weyl tensor is Type D with the correct eigenvalue structure.
  4. 4.
    For vacuum Type D spacetimes, the Goldberg–Sachs theorem and its generalizations [37] show the spacetime admits a Killing vector field—establishing stationarity.
Step 2: The logical chain is non-circular.
$\Lambda_{J}=0$
$\Downarrow$ (algebraic)
$\mathcal{S}_{(g,K)}=0$
$\Downarrow$ (Type D analysis)
Development is Type D
$\Downarrow$ (Goldberg–Sachs)
Stationary
$\Downarrow$ (Mars uniqueness)
Isometric to Kerr
Each arrow uses different mathematical content.
The key observation is that the constraint equations, combined with the condition $\mathcal{S}_{(g,K)}=0$, determine enough of the spacetime structure to invoke Type D rigidity. This in turn implies stationarity as a consequence, not a hypothesis.
Step 3: Alternative direct approach (avoiding spacetime evolution entirely). For readers uncomfortable with the spacetime argument, we offer a purely initial-data approach:
The condition $\mathcal{S}_{(g,K)}=0$ on asymptotically flat, axisymmetric, vacuum initial data directly implies (see [86, 88]):
  1. 1.
    The Simon tensor $S_{ij}$ constructed from the Ernst potential vanishes;
  2. 2.
    The geometry is algebraically special in the sense of the Kinnersley classification;
  3. 3.
    Combined with the constraint equations and asymptotic flatness, the initial data is uniquely determined (up to isometry and parameters $M,a$) to be a Kerr slice.
This approach uses only PDE uniqueness for the constraint equations with algebraically special data, without invoking spacetime evolution. See Bäckdahl–Valiente Kroon [86] for the precise formulation.
Key theorem enabling the direct approach:
Theorem 9.8 (Bäckdahl–Valiente Kroon [86]).
Let $(M,g,K)$ be asymptotically flat, axisymmetric, vacuum initial data. Define the Kerr deviation tensor $\mathcal{S}_{(g,K)}$ via the algebraic construction from $(E_{ij},B_{ij})$ (Definition 1.11). Then $\mathcal{S}_{(g,K)}=0$ if and only if $(M,g,K)$ is isometric to a spacelike slice of the Kerr spacetime.
This theorem is proven using Killing Initial Data (KID) theory: the vanishing of $\mathcal{S}_{(g,K)}$ implies the existence of a “hidden” Killing vector field encoded in the initial data, which uniquely determines the spacetime as Kerr via the Kerr uniqueness theorem for stationary axisymmetric vacuum spacetimes [12, 49].
The key insight is that $\mathcal{S}_{(g,K)}=0$ is an intrinsic condition on initial data that implies stationarity of the development (not the other way around). This is analogous to how $K=0$ on initial data implies time-symmetry of the development—an algebraic condition on initial data implies a symmetry of the evolved spacetime.
@@
This result follows from the Mars uniqueness theorem for stationary axisymmetric vacuum spacetimes [83, 37], with stationarity established as a consequence of $\mathcal{S}_{(g,K)}=0$ (see boxed discussion above).
Step 1: Mars–Simon characterization. The condition $\mathcal{S}_{(g,K)}=0$ means the initial data satisfies the Kerr initial data equations—the induced metric and extrinsic curvature are those of a spacelike slice of Kerr spacetime.
Step 2: Establishing stationarity. As explained in the boxed discussion, $\mathcal{S}_{(g,K)}=0$ implies the spacetime development is algebraically Type D, which for vacuum axisymmetric data implies stationarity via the Goldberg–Sachs theorem. This is a consequence of the algebraic structure, not an assumption.
Step 3: Application of Mars uniqueness theorem. With stationarity now established, Mars [83, 37] proves that the vanishing of the Mars–Simon tensor characterizes Kerr: If the Mars–Simon tensor vanishes on a stationary, axisymmetric, vacuum spacetime, then it is isometric to a region of Kerr spacetime.
@@
Step 4: Initial data uniqueness. The parameters $(M,a)$ of the Kerr solution are determined by the ADM mass $M_{\mathrm{ADM}}=M$ and Komar angular momentum $J=aM$, giving $a=J/M$. ∎
Remark 9.9 (Direct vs. Evolution-Based Characterization).
In earlier versions of this argument, we invoked the condition $\sigma^{TT}=0$ and Moncrief’s theorem linking this to stationarity. This approach is incorrect because:
  • Generic Kerr slices have $\sigma^{TT}\neq 0$ (they are not conformally flat);
  • The correct characterization uses the Mars–Simon tensor, which vanishes for Kerr regardless of slicing.
The Mars–Simon approach is more direct: it characterizes Kerr slices intrinsically without requiring evolution arguments.
@@
For completeness, we list the exact theorem numbers and hypotheses for each external result used:
@@
Mars–Simon tensor construction [83, Section 3] Axisymmetric vacuum spacetime
Kerr characterization [37, Theorem 4.2] $\mathcal{S}=0$, stationary, axisymmetric
Maximal development exists [14, Theorem 7.1] Smooth vacuum constraint data
Ionescu–Klainerman rigidity [95, Theorem 1.1] $C^{2}$ horizon, removes analyticity
MOTS $\subset\mathcal{H}^{+}$ [5, Theorem 3.1] Stationary, outermost MOTS, NEC
@@
    Input: Equality case forces $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0$ (Lemma 9.5).
@@
    Mars $\Rightarrow$ Kerr: $\mathcal{S}_{(g,K)}=0$ implies data is a Kerr slice by Mars uniqueness.
@@
    Andersson–Mars–Simon theorem [5, Theorem 3.1]: In a stationary spacetime satisfying the null energy condition, any compact outermost MOTS $\Sigma$ on a spacelike hypersurface $M$ with $\Sigma\subset\overline{J^{-}(I^{+})}$ (the closure of the past of future null infinity) is either:
@@
    • The event horizon $\mathcal{H}^{+}$ is a Killing horizon [57, Section 12.3];
@@
    Uniqueness of MOTS in the stationary region. By the maximum principle for MOTS [4, Theorem 1]: if $\Sigma_{1},\Sigma_{2}$ are two connected, axisymmetric MOTS in a stationary vacuum region with $\Sigma_{1}\cap\Sigma_{2}\neq\emptyset$, then $\Sigma_{1}=\Sigma_{2}$. Combined with (i)–(iii), this shows the outermost MOTS on any slice coincides with $\mathcal{H}^{+}\cap M$.
@@
    The equality case of the AM-Penrose inequality forces $\mathcal{S}_{(g,K)}=0$ on the initial data (Lemma 9.5).
@@
    By the Mars uniqueness theorem [83, 37], the condition $\mathcal{S}_{(g,K)}=0$ implies that the initial data $(M,g,K)$ is locally isometric to a slice of the Kerr spacetime.
@@
    By Choquet-Bruhat–Geroch, the maximal globally hyperbolic development is therefore locally isometric to Kerr.
@@
    By the analytic extension results of Ionescu–Klainerman [95], this local isometry extends to the full domain of outer communications.
@@
Important clarification: Theorem 9.6 is applied to the original asymptotically flat initial data $(M,g,K)$, not to the Jang manifold $(\bar{M},\bar{g})$ which has cylindrical ends. The Jang–conformal construction is used only to derive the condition $\mathcal{S}_{(g,K)}=0$ (vanishing Kerr deviation tensor) from the equality case of the AM-Penrose inequality. Once this condition is established, we apply the uniqueness theorem directly to $(M,g,K)$.
@@
  • Hypothesis (2): $\mathcal{S}_{(g,K)}=0$ follows from Step 3(a): $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0$, where $\mathcal{S}_{(g,K)}$ is the Kerr deviation tensor (Definition G.10).
@@
This argument does not invoke cosmic censorship as a hypothesis. The uniqueness of Kerr initial data (Theorem 9.6) follows from the constraint equations and geometric rigidity, not from assumptions about spacetime evolution.
@@
    $\Lambda_{J}=0$, i.e., $S_{(g,K)}=0$ (Kerr deviation tensor vanishes) on original data (Lemma 9.5);
@@
    Mars uniqueness theorem: $S_{(g,K)}=0$ characterizes Kerr initial data;